OpenStudy (anonymous):

There are 6 blue marbles,7 black marbles,4 orange marbles,and 3 green marbles in a bag. Once a marble is drawn,it is REPLACED. Find the probability of the outcome two orange marbles in a row a)2/25 b)1 c)1/25 d)2/5

OpenStudy (anonymous):

OK you tell me how many orange are there? How many total marbles are there?

OpenStudy (anonymous):

4 orange and total 20

OpenStudy (anonymous):

OK so on the first draw.. you put your hand in the bag and there are 4 orange out of the 20 = 4/20 reduced is 1/5 Then.. you put that orange marble back in the bag and you draw again. At that moment there are 4 orange in the bag out of 20 total marbles. So that time when you are it is still 4/20 which is 1/5 So the probablity = 1/5 * 1/5 = 1/25 So you would be able to draw an orange then an orange about 1 out of every 25 times.

OpenStudy (anonymous):

Thanks